### v4

@@ 5 Conclusion
In our work, we investigate the SSL for SAR ATR with our SAR-JEPA. SAR-JEPA combines two architectures to perform mask image modeling in gradient feature space for SAR images. By leveraging SAR domain features, we extend the applicability of MIM to SAR imagery, a type of noisy remote sensing data. Through extensive experimentation and analysis, we demonstrate the effectiveness of the proposed SAR-JEPA in obtaining high-quality representations for SAR ATR. Our research highlights the importance of embracing SSL to take advantage of the increasing availability of SAR data. The large pre-training samples. In the future, we will further extend the data and computational resources. give SSL the potential to achieve a foundation model for various downstream SAR ATR tasks across targets, scenes, and sensors.
However, it is worth noting that our pre-training dataset [^5] To the best of our knowledge, other SAR datasets can be further added to the pre-training, such as vehicle (Gotcha (Dungan et al., 2012), SIVED (Lin et al., 2023)), ship datasets (AIR-SARSHIP Sun et al. (2019), HRSID (Wei et al., 2020)), aircraft dataset (SADD Zhang et al. (2022)), and building datasets (FUSAR-Map Shi et al. (2021), OGSOD (Wang et al., 2023a)). However, how to further scale the SAR target data volume close to the million scale of ImageNet is a challenging problem. is not as extensive as ImageNet to fully exploit the potential of SSL for SAR ATR. In addition, it needs further investigation for other foundation model architectures, such as ConvNeXt. While our work primarily focuses on feature representation across different targets, it is essential to explore and discuss various tasks, including detection and segmentation. In the future, we plan to enhance our proposed scheme by improving data collection methods, refining model architectures, and exploring novel pre-training target features. Our ultimate goal is to successfully apply the foundation model with SSL in SAR target recognition.

### v5

@@ 5 Conclusion
In this study, we investigate SSL for SAR ATR using our SAR-JEPA. SAR-JEPA combines two architectures to perform mask image modeling in a gradient feature space for SAR images. By leveraging local patch masking and SAR domain features, we extend the applicability of MIM to SAR imagery, which is a type of noisy remote sensing data. Through extensive experimentation and analysis, we demonstrate the effectiveness of the proposed SAR-JEPA in obtaining high-quality representations for SAR ATR across various target categories and image conditions. This study highlights the importance of using SSL to exploit the increasing availability of SAR data. In the future, we will further extend the data and computational resources and provide SSL with the potential to achieve a foundation model for various downstream SAR ATR tasks across targets, scenes, and sensors.
However, it is noteworthy that our pretraining dataset needs to scale extensively, like ImageNet, to exploit the potential of SSL for SAR ATR. To the best of our knowledge, other SAR datasets can be further added to the pretraining, such as vehicle datasets Gotcha (Dungan et al., 2012), SIVED (Lin et al., 2023), ship datasets AIR-SARSHIP (Sun et al., 2019), HRSID (Wei et al., 2020), aircraft dataset SADD (Zhang et al., 2022), and building datasets FUSAR-Map (Shi et al., 2021), OGSOD (Wang et al., 2023a). However, scaling the SAR target data volume close to the million scale of ImageNet is a challenging problem. Another direction is to use the simulated data for SSL. For example, formula-driven supervised learning (Kataoka et al., 2022) does not use any natural images for SSL in computer vision and learns different fractal structures. In addition, more studies are required for other foundation model architectures, such as ConvNeXt. Although this study focuses primarily on feature representation across various targets, it is essential to explore and discuss various tasks, including detection and segmentation. In the future, we plan to enhance the proposed scheme by improving the data collection methods, refining the model architectures, and exploring novel pretraining target features. Our ultimate goal is to successfully apply the foundation model with SSL to SAR target recognition.